What is the volume of change of 2m3 of water at 20°C due to an applied pressure of 10 MPa?


1
Expert's answer
2021-10-19T15:13:09-0400

"\Delta P = 10MPa = 10^7Pa"

"V_0= 2m^3"

"K = 2.2*10^9 Pa\text{ bulk modulus of elasticity of water}"

"K= -V_0\frac{\Delta P}{\Delta V}"

"\Delta V= -V_0\frac{\Delta P}{K}=-2\frac{10^7}{2.2*10^9}=-0.91*10^{-2}m^3"


"\text{Answer: the volume of water decreased by }0.91*10^{-2}m^3"